OpenStudy (anonymous):

Perform the indicated operation and, if possible, simplify. Please show all of your work. -6/7 * 5 * (-14/27)

OpenStudy (accessdenied):

When we multiply fractions, we multiply the numerators together and the denominators together. \[ \frac{-6}{7} * \frac{5}{1} * \frac{-14}{27} = \frac{-6 * 5 * (-14)}{27 * 7}\] If we break the fraction down into its prime factorization, we can simplify because a/a =1 \[ \frac{-6*5*(-14)}{27 * 7} = \frac{-1 * 2 * 3 *5 *(-1)* 2 * 7}{3*3*3*7} \] -1 * -1 = 1, we can cancel out a 3/3 and 7/7. \[ \frac{2*5*2}{3*3} = \frac{20}{9} \]